# FedLPA: One-shot Federated Learning with Layer-Wise Posterior Aggregation

### v3

@@ 4.4 Ablation study
The hyper-parameter of our approach is $\lambda$ from Eq. 33, which controls variances of a priori normal distribution and guarantees $\mathbf{A}_{k}$ and $\mathbf{B}_{k}$ are positive semi-definite. In this part, we show results on FMNIST. All other Laplace Approximations are sensitive to the hyper-parameter $\lambda$ based on their experimental results, Table 4 shows that our approach is relatively robust. Based on our numerical results, we set $\lambda=0.001$ by default for our method FedLPA.
We also conduct the experiments when the local epochs are 10,20,50,100. More experiments are available in Appendix M.3, which shows that our methods outperform all the baselines in all kinds of scenarios without requiring extensive tuning.
@@ 4.5 Communication and computation overhead
We conduct experiments on CIFAR-10 on a single 2080Ti GPU to estimate the overall communication and computation overhead. We set the number of clients is 10. Table 4 shows the numerical results on FedLPA and baselines. Details of the overhead evaluation are referred to Appendix M.8 and M.9. Our observations reveal that FedLPA is slightly slower than FedNova, SCAFFOLD, FedAvg, and FedProx, while much faster than DENSE. FedLPA also has significantly improved the one-shot learning performance of the above four approaches. Similarly, FedLPA performs moderately incremental communication overhead while outperforming other baseline approaches on learning performance, as one-shot FL introduces heavy computation overhead while communication overhead is usually small. It is noteworthy that FedLPA strikes a favorable balance between computation and communication overhead, making it the most promising approach for one-shot FL.
@@ 4.6 Extension to multiple rounds
We conduct experiments on MNIST with 10 clients and data partitioning $p_{k}$ Dir($\beta=0.5$). The results are shown in Figure 2. As DENSE could not support multiple rounds, we compare our methods with FedAvg, FedNova, SCAFFOLD, and FedProx. FedLPA achieves the highest accuracy in the first round, denoting the strongest learning capabilities in a one-shot setting. With the increment in the number of rounds, the performances of FedLPA increase slower than the other baseline approaches. This figure shows that the joint approach (ours (one round) then FedAvg) that utilizes FedLPA in the first round and then adopts other baseline methods may be most promising to save communication and computation resources in the multiple-round federated learning scenario. More details are in Appendix G.1.
@@ 4.7 Supplementary experiments
Experiments for privacy concerns, experiments on different local epoch numbers, experiments in extreme settings (the number of clients=5, $\beta=0.001$), experiments for aggregation visualization, experiments with more methods, experiments with more complex network structures, experiments with more complex datasets, ablation experiments analyzing the number of approximation iterations of FedLPA, experiments with different proportion of data samples can be found in Appendix.
@@ 5 Conclusions
In this work, we design a novel one-shot FL algorithm FedLPA to better model the global parameters in effective one-shot federated learning. We propose a method that could aggregate the local clients in a layer-wise manner with their posterior approximation via block-diagonal empirical Fisher information matrices, which could effectively capture the accurate statistics of a locally biased dataset. Overall, FedLPA stands out as the most practical and efficient framework that conducts data-free one-shot FL, particularly well-suited for high data heterogeneity in various settings, considering it significantly outperforms other baselines with extensive experiments.
@@ Appendix A The FedLPA algorithm
The proposed algorithm follows the same paradigm as the standard one-shot federated learning framework. Each client follows the local training procedure as Algorithm 1. The global aggregation is illustrated in Algorithm 2.
@@
Then, We give a concrete example to show the dimensions of different matrices using a fully-connected neural network model with architecture 784-256-64-10 as in Appendix M. Then, the $\mathbf{M}_{k_{1}}$ is 784x256+256, $\mathbf{M}_{k_{2}}$ is 256x64+64, $\mathbf{M}_{k_{3}}$ is 64x10+10. The $\mathbf{A}_{k_{1}}$ is 785x785, $\mathbf{A}_{k_{2}}$ is 257x257, $\mathbf{A}_{k_{3}}$ is 65x65. The $\mathbf{B}_{k_{1}}$ is 256x256, $\mathbf{B}_{k_{2}}$ is 64x64, $\mathbf{B}_{k_{3}}$ is 10x10. Then the $\mathbf{F}_{k_{1}}$ is (785x785)x(256x256), $\mathbf{F}_{k_{2}}$ is (257x257)x(64x64), $\mathbf{F}_{k_{3}}$ is (65x65)x(10x10). The $\mathbf{F}_{k}$ is (785x785+257x257+65x65)x(256x256+64x64+10x10).
@@
Algorithm 1 FedLPA Local Training
1:
2: Training set $\mathcal{D}_{k}$ of the $k$-th. client
3: Batch size $C$
4: The number of layers $L$
5: Factor $\lambda$ and $\{\pi_{1},\pi_{2},\ldots,\pi_{L}\}$
6: The global parameters $\{\bar{\mathbf{M}}_{1},\bar{\mathbf{M}}_{2},\ldots,\bar{\mathbf{M}}_{L}\}$
7:
8:Initialize weights $\{\mathbf{W}_{k_{1}},\mathbf{W}_{k_{2}},\ldots,\mathbf{W}_{k_{L}}\}$ using $\{\bar{\mathbf{M}}_{1},\bar{\mathbf{M}}_{2},\ldots,\bar{\mathbf{M}}_{L}\}$
9:Obtain optimal weights $\{\mathbf{W}^{*}_{k_{1}},\mathbf{W}^{*}_{k_{2}},\ldots,\mathbf{W}^{*}_{k_{L}}\}$leftarrow Train the model using gradient descent $\triangleright$ Optimization
10:
11:for all $c$ from 1 to $C$ do $\triangleright$ Iterate batch data
12:     Initialize $\mathbf{a}_{k_{0}}^{(c)}=\mathbf{x}^{(c)}\in\mathcal{D}_{k}$
13:     for all $l$ from 1 to $L$ do $\triangleright$ Forward pass
14:         $\hat{\mathbf{a}}_{k_{l-1}}^{(c)}$leftarrow coordinate-wised append value 1 to $\mathbf{a}_{k_{l}-1}^{(c)}$
15:         $\mathbf{h}_{k_{l}}\leftarrow\mathbf{W}^{*}_{k_{l}}\hat{\mathbf{a}}_{k_{l-1}}^{(c)}$
16:         $\mathbf{a}_{k_{l}}^{(c)}\leftarrow\sigma_{k_{l}}\left(\mathbf{h}_{k_{l}}\right)$
17:     end for
18:     $\nabla\mathbf{a}_{k_{L}}^{(c)}\leftarrow\frac{\partial L(\mathbf{y}^{(c)},\mathbf{z})}{\partial\mathbf{z}}|_{\mathbf{z}=\mathbf{a}_{k_{L}}^{(c)}}$triangleright Loss derivative computation
19:     for all $l$ from $L$ down to 1 do $\triangleright$ Backwards pass
20:         $\mathbf{g}_{k_{l}}^{(c)}\leftarrow\nabla\mathbf{a}_{k_{l}}^{(c)}\odot\sigma_{k_{l}}^{\prime}\left(\mathbf{h}_{k_{l}}\right)$
21:         $\hat{\mathbf{b}}_{k_{l}}\leftarrow\mathbf{g}_{k_{l}}^{(c)}\hat{\mathbf{a}}_{k_{l-1}}^{(c)\top}$
22:         $\nabla\mathbf{a}_{k_{l-1}}^{(c)}\leftarrow\mathbf{W}^{*\top}_{k_{l}}\mathbf{g}_{k_{l}}^{(c)}$
23:         $\mathbf{A}_{k_{l}}^{(c)}\leftarrow\hat{\mathbf{a}}_{k_{l-1}}^{(c)}\hat{\mathbf{a}}_{k_{l-1}}^{(c)\top}+\pi_{l}\sqrt{\lambda}\mathbf{I}$
24:         $\mathbf{B}_{k_{l}}^{(c)}\leftarrow\hat{\mathbf{b}}_{k_{l}}^{(c)}\hat{\mathbf{b}}_{k_{l}}^{(c)\top}+\frac{1}{\pi_{l}}\sqrt{\lambda}\mathbf{I}$
25:         Compute running averages of $\mathbf{A}_{k_{l}}$ and $\mathbf{B}_{k_{l}}$ over batches
26:     end for
27:end for
28:for all $l$ from 1 to $L$ do
29:     $\mathbf{M}_{k_{l}}\leftarrow\mathbf{W}^{*}_{k_{l}}$
30:end for
31:return $\{\mathbf{M}_{k_{l}},\mathbf{A}_{k_{l}},\mathbf{B}_{k_{l}}|l=1,...,L\}$
Algorithm 2 FedLPA Global Aggregation
1:Input: clients $K$, layers $L$
2:Initialize global weight $\bar{\mathbf{W}}_{l}$ of layer $l=1,...,L$
3:clients executes:
4:Initialize local model
5:for k = 1, …, K do
7:end for
8:Server executes:
9:for l = 1, …, L do
10:     $\bar{\mathbf{A}}_{l}\leftarrow\sum_{k}^{K}\mathbf{A}_{k_{l}}$
11:     $\bar{\mathbf{B}}_{l}\leftarrow\sum_{k}^{K}\mathbf{B}_{k_{l}}$
14:     $\bar{\mathbf{W}}_{l}\leftarrow\bar{\mathbf{M}}_{l}$
15:end for
Appendix B The standard Laplace approximation
Here we give the details of Eq.2:
By taking the Bayesian equation into the global posterior, the logarithm of the global posterior is the average of the logarithm of local posteriors.
$\log p(\bm{\theta}|\mathcal{D})=\sum_{s\in\mathcal{D}}\log p(s|\bm{\theta})+\log p(\bm{\theta})-\sum_{s\in\mathcal{D}}\log p(s)$
$=\sum_{k=1}^{K}\log p(\mathcal{D}_{k}|\bm{\theta})-\sum_{k=1}^{K}\log p(\mathcal{D}_{k})+\log p(\bm{\theta})$
$\propto\sum_{k=1}^{K}\log p(\mathcal{D}_{k}|\bm{\theta})+\sum_{k=1}^{K}\frac{1}{k}\log p(\bm{\theta})-\sum_{k=1}^{K}\log p(\mathcal{D}_{k})$
$\propto\sum_{k=1}^{K}\log p(\bm{\theta}|\mathcal{D}_{k})$
That is how we derive $\sum_{k=1}^{K}\log p(\mathcal{D}_{k}|\bm{\theta})\propto\sum_{k=1}^{K}\log p(\bm{\theta}|\mathcal{D}_{k})$.
Thus we have $\prod_{k=1}^{K}p(\mathcal{D}_{k}|\bm{\theta})\propto\prod_{k=1}^{K}p(\bm{\theta}|\mathcal{D}_{k})$.
By doing so, we could model the $p(\bm{\theta}|\mathcal{D})$ via $\prod_{k=1}^{K}p(\bm{\theta}|\mathcal{D}_{k})$ instead of $\prod_{k=1}^{K}p(\mathcal{D}_{k}|\bm{\theta})$ to introduce our method.
Then, the standard Laplace approximation [43] begins by taking the second-order Taylor expansion around an optimum of posterior:
$\log p(\bm{\theta}|\mathcal{D})\approx\log p\left(\bm{\theta}^{*}|\mathcal{D}\right)-\frac{1}{2}\left(\bm{\theta}-\bm{\theta}^{*}\right)^{\top}\bar{\mathbf{H}}\left(\bm{\theta}-\bm{\theta}^{*}\right),$ (15)
where $\bm{\theta}^{*}$ is the optimal parameter vector, $\bar{\mathbf{H}}=\mathbb{E}_{s\in\mathcal{D}}[\mathbf{H}]$ is the average Hessian of the negative log posterior over a dataset $\mathcal{D}$.
Eq. 5 ignores the first-order term because the first-order derivative is almost zero near the optimal. Terms of a higher order than the second-order terms are ignored because the curvature near the optimal value is assumed to be quadratic. Given the optimal $\bm{\theta^{*}}$ and dataset $\mathcal{D}$, $\log p\left(\bm{\theta^{*}}|\mathcal{D}\right)$ is a determined constant, so that the exponential format of Eq. 5 is proportional to a multi-variates Gaussian distribution:
$\mathcal{N}(\bm{\theta^{*}},\bar{\mathbf{H}}^{-1})\propto\exp\{\log p\left(\bm{\theta^{*}}|\mathcal{D}\right)-\frac{1}{2}\left(\bm{\theta}-\bm{\theta^{*}}\right)^{\top}\bar{\mathbf{H}}\left(\bm{\theta}-\bm{\theta^{*}}\right)\}.$ (16)
Therefore, it is reasonable to approximate a posterior as a Gaussian distribution with expectations $\mathbf{\mu}=\bm{\theta}^{*}$ and co-variances $\mathbf{\Sigma}=\bar{\mathbf{H}}^{-1}$.
$p(\bm{\theta}|\mathcal{D})\equiv\bm{\theta}\sim\mathcal{N}(\mathbf{\mu},\mathbf{\Sigma}).$ (17)
With a generally, we regard the prior $p(\bm{\theta})\equiv\bm{\theta}\sim\mathcal{N}(0,\lambda^{-1}\mathbf{I})$ as a Normal distribution.
Appendix C Further discussion of Eq. 5
Eq. 5 ignores the first-order term because the first-order derivative is almost zero near the optimal. Terms of a higher order than the second-order terms are ignored because they are much less significant and the curvature near the optima is assumed to be quadratic [43].
In fact, the Gaussian distribution is employed to approximate the small neighborhood of the posteriors based on the MAP estimate. This choice is made because the empirical Fisher information matrix we utilize in the end is equivalent to the expectation of the Hessian of the negative log posterior, as detailed in Appendix E. We are not approximating the entire posterior as Gaussian but rather approximating the local optimization curve near the posterior parameters as Gaussian, where the distance of the points on the curve to the true posterior point is sufficiently small. Thus, it is reasonable to approximate global and local posteriors as multi-variates Gaussian distributions with expectations $\bar{\bm{\mu}}=\bm{\theta}^{*}$ and $\bm{\mu_{k}}=\bm{\theta}_{k}^{*}$; co-variances $\bar{\mathbf{\Sigma}}=\bar{\mathbf{H}}^{-1}$ and $\mathbf{\Sigma}_{k}=\bar{\mathbf{H}}_{k}^{-1}$ [44].
Appendix D Futher discussion of Eq. 7
Combining Eq. 2 and Eq. 6, and taking the logarithm of both sides, we obtain: $(\bm{\theta}-\bar{\bm{\mu}})^{T}\bar{\mathbf{\Sigma}}^{-1}(\bm{\theta}-\bar{\bm{\mu}})=\sum_{k}^{K}{(\bm{\theta}-\bar{\bm{\mu}})^{T}\mathbf{\Sigma}^{-1}_{k}(\bm{\theta}-\bar{\bm{\mu}})}+const$
Differentiating the above equation with respect to $\bm{\theta}$ yields, $\bar{\mathbf{\Sigma}}^{-1}(\bm{\theta}-\bar{\bm{\mu}})=\sum_{k}^{K}{\mathbf{\Sigma}^{-1}_{k}(\bm{\theta}-\bm{\mu_{k}})}$.
Next, differentiating the first-order derivative formula with respect to $\bm{\theta}$ again, which is the second derivative of the original equation, we get $\mathbf{\Sigma}^{-1}=\sum_{k}^{K}\mathbf{\Sigma}^{-1}_{k}$. Substituting this $\mathbf{\Sigma}^{-1}$ result back into the first-order derivative result yields $\bar{\bm{\mu}}$. Thus, we get Eq. 7.
Appendix E Estimate Hessian matrix using empirical Fisher information matrix
We can find that the co-variances are a sum of prior co-variances and the expectations of the second derivation of likelihood.
$\mathbf{\Sigma}^{-1}$ $=\bar{\mathbf{H}}$ (18)
$=-\nabla^{2}\log p(\bm{\theta}|\mathcal{D})$
=$=-\{\nabla^{2}\log p(\mathcal{D}|\bm{\theta})+\nabla^{2}\log p(\bm{\theta})-\nabla^{2}p(\mathcal{D})\}$
$=-\mathbb{E}_{s\in\mathcal{D}}\left[\nabla^{2}\log p(s|\bm{\theta})\right]+\lambda\mathbf{I}-0$
$=-\frac{1}{|\mathcal{D}|}\sum_{i=1}^{|\mathcal{D}|}\nabla^{2}\log p(s_{i}|\bm{\theta})+\lambda\mathbf{I}$
In the papers [48, 62, 50, 17, 51] and our method, we all assume that $p(\theta)$ follows the normal distribution, which is a practical assumption and is shown in Eq. 17 in our paper.
However, $\bar{\mathbf{H}}$ is intractable to compute and store due to many parameters in modern neural networks. An efficient method is to approximate $\bar{\mathbf{H}}$ as a diagonal matrix using the empirical Fisher information matrix $\tilde{\mathbf{F}}$ [47].
For a neural network, a $\tilde{\mathbf{F}}$ is defined as below:
$\tilde{\mathbf{F}}=\mathbb{E}_{p(s|\bm{\theta})}\left[\nabla\log p(s|\bm{\theta})\nabla\log p(s|\bm{\theta})^{\top}\right]$ (19)
where $p(s|\bm{\theta})$ is the likelihood on data point $x$.
The $\tilde{\mathbf{F}}$ has an obvious property that it is equivalent to the expectation of the Hessian of the negative log posterior, as shown below:
$\bar{\mathbf{H}}$ $=\mathbb{E}_{p(s|\bm{\theta})}\left[\mathbf{H}_{-\log p(s|\bm{\theta})}\right]$ (20)
$=-\mathbb{E}_{p(s|\bm{\theta})}\left[\mathbf{H}_{\log p(s|\bm{\theta})}\right]$
$=-\mathbb{E}_{p(s|\bm{\theta})}\left[\mathbf{J}\left(\frac{\nabla p(s|\bm{\theta})}{p(s|\bm{\theta})}\right)\right]$
$=-\mathbb{E}_{p(s|\bm{\theta})}\left[\frac{\mathbf{H}_{p(s|\bm{\theta})}p(s|\bm{\theta})-\nabla p(s|\bm{\theta})\nabla p(s|\bm{\theta})^{\top}}{p(s|\bm{\theta})p(s|\bm{\theta})}\right]$
$=-\mathbb{E}_{p(s|\bm{\theta})}\left[\frac{\mathbf{H}_{p(s|\bm{\theta})}p(s|\bm{\theta})}{p(s|\bm{\theta})p(s|\bm{\theta})}-\frac{\nabla p(s|\bm{\theta})\nabla p(s|\bm{\theta})^{\top}}{p(s|\bm{\theta})p(s|\bm{\theta})}\right]$
$=-\mathbb{E}_{p(s|\bm{\theta})}\left[\frac{\mathbf{H}_{p(s|\bm{\theta})}}{p(s|\bm{\theta})}-\left(\frac{\nabla p(s|\bm{\theta})}{p(s|\bm{\theta})}\right)\left(\frac{\nabla p(s|\bm{\theta})}{p(s|\bm{\theta})}\right)^{\top}\right]$
$=-\mathbb{E}_{p(s|\bm{\theta})}\left[\frac{\mathbf{H}_{p(s|\bm{\theta})}}{p(s|\bm{\theta})}\right]+\mathbb{E}_{p(s|\bm{\theta})}\left[\left(\frac{\nabla p(s|\bm{\theta})}{p(s|\bm{\theta})}\right)\left(\frac{\nabla p(s|\bm{\theta})}{p(s|\bm{\theta})}\right)^{\top}\right]$
$=\mathbb{E}_{p(s|\bm{\theta})}\left[\nabla\log p(s|\bm{\theta})\nabla\log p(s|\bm{\theta})^{\top}\right]-\int\frac{\mathbf{H}_{p(s|\bm{\theta})}}{p(s|\bm{\theta})}p(s|\bm{\theta})\mathrm{d}x$
$=\tilde{\mathbf{F}}-\mathbf{H}_{\int p(s|\bm{\theta})\mathrm{d}x}$
$=\tilde{\mathbf{F}}-\mathbf{H}_{1}$
$=\tilde{\mathbf{F}}$
However, the expectation of the likelihood $p(s|\bm{\theta})$ is usually intractable to compute. Therefore, a common method is to use an empirical distribution $q(s)\approx p(s|\bm{\theta})$ to approximate the likelihood, and $q(s)$ is always assumed to be $p(\mathcal{D})$.
$\tilde{\mathbf{F}}$ $\approx\mathbb{E}_{q(s)}\left[\nabla\log p(s|\bm{\theta})\nabla\log p(s|\bm{\theta})^{\top}\right]$ (21)
$=\mathbb{E}\left[\nabla\log p(s|\bm{\theta})\nabla\log p(s|\bm{\theta})^{\top}\right]$
$=\frac{1}{|\mathcal{D}|}\sum_{i=1}^{|\mathcal{D}|}\nabla\log p(s_{i}|\bm{\theta})\nabla\log p(s_{i}|\bm{\theta})^{\top}$
Eq. 21 is called the empirical Fisher information matrix, which is computable. Therefore, a diagonal matrix of $\bar{\mathbf{H}}$ can be estimated by the diagonal of empirical $\tilde{\mathbf{F}}$. We denote $\mathrm{diag}(\cdot)$ as the diagonal of a matrix.
$\mathrm{diag}(\bar{\mathbf{H}})$ $=\mathrm{diag}(\tilde{\mathbf{F}}+\lambda\mathbf{I})$ (22)
$\approx\frac{1}{|\mathcal{D}|}\sum_{i=1}^{|\mathcal{D}|}\mathrm{diag}(\nabla\log p(x|\bm{\theta})\nabla\log p(x|\bm{\theta})^{\top})+\lambda\mathbf{I}$
Appendix F Comparison with the previous methods
To the best of our knowledge, we are the first to consider the posterior inference problem in one-shot scenario. Note that the approach [16] requires a lengthy burn-in period before conducting posterior inference, for instance, 400 rounds, and it updates global model parameters by modifying the covariance-aggregated local models. It means that the algorithm [16] necessarily requires multiple iterations and cannot be used in a one-shot scenario. In contrast, our method FedLPA only requires immediate variational inference after training the local model, ensuring higher flexibility and efficiency in one-shot scenario.

### v5

@@ 4.3 Scalability
Table 3: Experiments with different proportions of data samples.
100% 55.33$\pm$0.06 68.20$\pm$0.04 73.33$\pm$0.06
80% 53.88$\pm$1.14 65.47$\pm$0.02 73.17$\pm$0.05
60% 53.15$\pm$0.82 64.80$\pm$0.71 72.40$\pm$0.29
40% 53.20$\pm$0.21 64.10$\pm$0.40 70.02$\pm$0.17
20% 45.71$\pm$0.13 62.15$\pm$0.03 68.54$\pm$2.02
4.4 Experiments with different proportions of data samples
We have added the experiments with our method on the same experiment setting with 10 clients. We conducted experiments on FMNIST datasets with $\beta$=0.1, 0.3 and 0.5. The performance changes w.r.t the number of data samples are shown in Table 3. We could see that our method FedLPA could yield satisfactory results even with only 20% data samples under multiple settings.
@@ 4.5 Ablation study
The hyper-parameter of our approach is $\lambda$, which controls variances of a priori normal distribution and guarantees $\mathbf{A}_{k}$ and $\mathbf{B}_{k}$ are positive semi-definite. In this part, we show results on FMNIST. All other Laplace Approximations are sensitive to the hyper-parameter $\lambda$ based on their experimental results, Table 5 shows that our approach is relatively robust. Based on our numerical results, we set $\lambda=0.001$ by default for our method FedLPA.
We also conduct the experiments when the local epochs are 10,20,50,100. More experiments are available in Appendix G.2, which shows that our methods outperform all the baselines in all kinds of scenarios without requiring extensive tuning.
@@ 4.6 Communication and computation overhead
We conduct experiments on CIFAR-10 on a single 2080Ti GPU to estimate the overall communication and computation overhead. We set the number of clients is 10. Table 5 shows the numerical results on FedLPA and baselines. Details of the overhead evaluation are referred to Appendix G.6 and G.7. Our observations reveal that FedLPA is slightly slower than FedNova, SCAFFOLD, FedAvg, and FedProx, while much faster than DENSE. FedLPA also has significantly improved the one-shot learning performance of the above four approaches. Similarly, FedLPA performs moderately incremental communication overhead while outperforming other baseline approaches on learning performance, as one-shot FL introduces heavy computation overhead while communication overhead is usually small. It is noteworthy that FedLPA strikes a favorable balance between computation and communication overhead, making it the most promising approach for one-shot FL.
@@ 4.7 Extension to multiple rounds
We conduct experiments on MNIST with 10 clients and data partitioning $p_{k}$ Dir($\beta=0.5$). The results are shown in Figure 2. As DENSE could not support multiple rounds, we compare our methods with FedAvg, FedNova, SCAFFOLD, and FedProx. FedLPA achieves the highest accuracy in the first round, denoting the strongest learning capabilities in a one-shot setting. With the increment in the number of rounds, the performances of FedLPA increase slower than the other baseline approaches. This figure shows that the joint approach (ours (one round) then FedAvg) that utilizes FedLPA in the first round and then adopts other baseline methods may be most promising to save communication and computation resources in the multiple-round federated learning scenario.
@@ 4.8 Supplementary experiments
Experiments for privacy concerns, experiments on different local epoch numbers, experiments in extreme settings (the number of clients=5, $\beta=0.001$), experiments with more methods, experiments with more complex network structures, experiments with more complex datasets, ablation experiments analyzing the number of approximation iterations of FedLPA can be found in Appendix.
@@ 5 Conclusions
In this work, we design a novel one-shot FL algorithm FedLPA to better model the global parameters in effective one-shot federated learning. We propose a method that could aggregate the local clients in a layer-wise manner with their posterior approximation via block-diagonal empirical Fisher information matrices, which could effectively capture the accurate statistics of a locally biased dataset. Overall, FedLPA stands out as the most practical and efficient framework that conducts data-free one-shot FL, particularly well-suited for high data heterogeneity in various settings, considering it significantly outperforms other baselines with extensive experiments. Our FedLPA is available in https://github.com/lebronlambert/FedLPA_NeurIPS2024.
Acknowledgments and Disclosure of Funding
We thank the anonymous reviewers. The authors would like to thank Yiqun Diao from National Univeristy of Singapore for his valuable comments to improve this work. Dr. Jialin Li is supported by the Singapore Ministry of Education Academic Research Fund Tier 1 (T1 251RES2104) and Tier 2 (MOE-T2EP20222-0016).
@@ Appendix A The FedLPA algorithm
The proposed algorithm follows the same paradigm as the standard one-shot federated learning framework. Each client follows the local training procedure as shown in the paper. The global aggregation is illustrated in Algorithm 1.
@@
Then, we give a concrete example to show the dimensions of different matrices using a fully-connected neural network model with architecture 784-256-64-10 as in Appendix G. Then, the $\mathbf{M}_{k_{1}}$ is 784x256+256, $\mathbf{M}_{k_{2}}$ is 256x64+64, $\mathbf{M}_{k_{3}}$ is 64x10+10. The $\mathbf{A}_{k_{1}}$ is 785x785, $\mathbf{A}_{k_{2}}$ is 257x257, $\mathbf{A}_{k_{3}}$ is 65x65. The $\mathbf{B}_{k_{1}}$ is 256x256, $\mathbf{B}_{k_{2}}$ is 64x64, $\mathbf{B}_{k_{3}}$ is 10x10. Then the $\mathbf{F}_{k_{1}}$ is (785x785)x(256x256), $\mathbf{F}_{k_{2}}$ is (257x257)x(64x64), $\mathbf{F}_{k_{3}}$ is (65x65)x(10x10). The $\mathbf{F}_{k}$ is (785x785+257x257+65x65)x(256x256+64x64+10x10).
@@
Appendix B Comparison with the previous methods
To the best of our knowledge, we are the first to consider the posterior inference problem in the one-shot scenario. Note that the approach [16] requires a lengthy burn-in period before conducting posterior inference, for instance, 400 rounds, and it updates global model parameters by modifying the covariance-aggregated local models. It means that the algorithm [16] necessarily requires multiple iterations and cannot be used in a one-shot scenario. In contrast, our method FedLPA only requires immediate variational inference after training the local model, ensuring higher flexibility and efficiency in the one-shot scenario.